OpenStudy (anonymous):

How many distinct factors does the number 144 have?

OpenStudy (anonymous):

Start dividing by the easiest numbers first. Start with 2 and keep dividing as long as you keep getting an integer as your answer. Then try 3, 5, 7 until you are at a prime integer remainder.

OpenStudy (anonymous):

you could prime factor 144:\[144=12^2=4^2\cdot 3^2=2^4\cdot 3^2\]If you have a factor of 144, its got to have some 2s and/or 3s in it. It could have {0,1,2,3,4} 2s and {0,1,2} 3s. So you have 5 choices for the power of 2, and 3 choices for the power of 3. So there are 5*3=15 total factors of 144.

OpenStudy (anonymous):

i think there is 15 but i could be wrong

OpenStudy (anonymous):

oh thank you:)